# Discrete Max-Focusing

### v2

@@ 3.3 Special Case: Compressible States
We stress that in such settings, our novel definition of future-nonexpanding, Def. 28, and our Discrete Max-Focusing Conjecture, Conj. 32, are both still needed. Their purpose is to deal with edge points where the classical expansion is not well-defined, and this problem is unrelated to the quantum corrections to the area term.
@@ 3.4 Special Case: Numerical Max-Expansion
Let $a$ be a wedge whose edge ð$\eth a$ is acausal and $C^{2}$-smooth. (That is, no two points in ð$\eth a$ are causally connected; and ð$\eth a$ is a $C^{2}$ embedded submanifold.) Let $y$ collectively denote the transverse coordinates that label points in ð$\eth a$, and let $h$ be the determinant of the induced metric on ð$\eth a$. Let $V:\eth a\to[0,1]$, $y\to V(y)$ be a piecewise continuous function. Let $a^{+}_{V,\lambda}$ be the outward deformation of $a$ along the unique null generator of $H^{+}(a^{\prime})$ that originates at each point $y$ in $O$, by an affine distance $\lambda V(y)$. (Note that $a^{+}_{V,\lambda}$ exists for all sufficiently small $\lambda$.) Then there exists a continuous function $\Theta_{\rm max}^{+}:O\to\mathbb{R}$ such that
@@
The above definition can be generalized somewhat, though we will not try to do this carefully here: if ð$\eth a$ is not acausal or not smooth (or neither), $\Theta^{\pm}(a,p)$ can still be defined at any point $p\in\eth a$ that is acausal to the rest of ð$\eth a$ and possesses an open neighborhood $O$ such that ð$\eth a\cap O$ is a smooth embedded submanifold. Moreover, even at a point $p$ where ð$\eth a$ is not $C^{2}$, one of the expansions may exist: namely, if one of $H^{\pm}(a^{\prime})$ is smooth in a neighborhood of $p$.
@@
Since $a$ is future-nonexpanding at $p$, there exists an open neighborhood $O$ of $p$ such that any future-outward deformation $b$ along $H^{+}(a)$ that remains within $O$ obeys $H_{\rm max,gen}(b|a)\leq 0$. Let $V$ and $a^{+}_{V,\lambda}$ be defined as in Def. 39. Since $H_{\rm max,gen}(a^{+}_{V,\lambda}|a)$ is nonpositive for sufficiently small $\lambda$, and since Eq. (30) holds for all nonpositive functions $V:O\to\mathbb{R}_{0}^{+}$, one finds that $\Theta_{\rm max}^{+}\leq 0$. ∎
@@
By Def. 39, there exists an open neighborhood $O$ of $p$ such that $\Theta_{\rm max}^{+}(a;q)<0$ for all $q\in O$. We may now choose any piecewise continuous function $V:0\to[0,1]$, giving rise to a family of deformations $a^{+}_{V,\lambda}$ as described in Def. 39. Since $\Theta_{\rm max}^{+}(a;p)$ is strictly negative, by continuity, $H_{\rm max,gen}(a^{+}_{V,\lambda}|a)$ must be negative for all sufficiently small $\lambda$. After restricting $O$ to such values for all possible choices of $V$, Eq. (30) implies that Def. 28 is met. ∎
@@ 4.1 No Geodesics Enter Lightsheets
Discrete Max-Focusing would be violated already at the level of differential geometry if it was possible for new null geodesics to enter $L^{+}(a)$, i.e., if we could choose ð$\eth c$ to intersect null geodesics that were not present on ð$\eth b$. Then by choosing ð$\eth c$ to include only a small neighborhood of these geodesics and otherwise agree with ð$\eth b$, the area contribution to $H_{\rm max,gen}(c|b)$ would be expected to dominate and would be positive.
@@ 4.2 Classical Focusing
By Def. 28, $a$ is future-nonexpanding at any point $p$ where ð$\eth a$ is differentiable, if and only if the expansion $\theta^{+}$ of the null generators of $H^{+}(a^{\prime})$ is nonpositive at $a$. Then both Theorem 34 and Conjecture 32 follow immediately from the Raychaudhuri equation Wald:1984rg ,
@@ 4.3 Discrete Max-Focusing Implies the QNEC
Let $a$ be a wedge in Minkowski space (with $G=0$) whose edge ð$\eth a$ is acausal and smooth. Let $p\in\eth a$, and let the vector $k^{\mu}$ be null, future and outward-directed, and orthogonal to ð$\eth a$ at $p$. We require conditions on the extrinsic curvature of ð$\eth a$ at $p$ that are necessary for the QNEC to be regularization-scheme independent Akers:2017ttv . In particular, the classical null expansion and shear must vanish at $p$: $\theta^{+}=\varsigma^{+}_{\mu\nu}=0$. Let $\psi$ be the state of the quantum fields on $M$ (pure or mixed). Now consider $M$ with $n$ copies of the same quantum field theory, in the state $\psi^{\otimes n}$, and turn on $G\neq 0$. Our argument will involve taking the limit as $G$ is turned back off. Therefore, we can take $G$ to be arbitrarily small at the present step.
@@
Neither the identifications of $a(G,n)$ and $p(G,n)$, nor the deformation into $b(G,n)$ are unique, even while satisfying the criteria of Ref. Akers:2017ttv . The only important requirement is that they be chosen such that $p(G,n)\to p$, $a(G,n)\to a$, and $b(G,n)\to a$ as $G\to 0$ at fixed $n$ [^4] To define an appropriate notion of continuity, one can represent the different spacetimes $M(G,n)$ as Minkowski space with a perturbative gravitational field. The latter property can be satisfied because $\Theta_{\rm max}^{+}[b(G,n);p(G,n)]-\theta^{+}[a(G,n);p(G,n)]\sim O(G)$.
We may now apply Discrete Max-Focusing to $b(G,n)$. Since $\Theta_{\rm max}^{+}[b(G,n);p(G,n)]<0$ and $b(G,n)$ can be chosen smooth, there exists an open neighborhood $O$ of $p(G,n)$ such that $\Theta_{\rm max}^{+}[b(G,n);q]<0$ for all $q\in\eth b(G,n)\cap O$; hence ð$\eth b(G,n)\cap O\subset\eth b(G,n)^{+}$. Let $c(G,n)\supset b(G,n)$ be an outward deformation of $b(G,n)$ along $H^{+}[b^{\prime}(G,n)]$ such that ð$\eth b(G,n)\setminus O=\eth c(G,n)\setminus O$. Let $q\in\eth c$ lie on the future null generator originating at $p$. We may choose $c(G,n)$ such that in the limit as $G\to 0$ at fixed $n$, it approaches a fixed wedge $c$ whose edge lies on $H^{+}(a^{\prime})$. (Recall that $a(G,n)\to a$, and $b(G,n)\to a$ as $G\to 0$ at fixed $n$.) Theorem 34 implies
@@
Here the quantum part is denoted with superscript $Q$ and defined as a functional derivative analogously to Eq. (30) (see Def. 39 for details):
@@
[…]
@@
[…]
@@ 5 Discrete-Max-Nonexpanding Wedges and Their Properties
$k$ is past- and future-nonexpanding at $p\in\eth k$ $k$ is antinormal at $p$ $k$ is accessible from $a$ [^6] This terminology was introduced in Ref. Akers:2023fqr for the special case where $a$ lies on the conformal boundary of Anti-de Sitter space.
$k$ is future- (past-) nonexpanding but no past (future) outward deformation of $k$ is future- (past-) nonexpanding $k$ is future-marginal (past-marginal) [^7] In the classical limit and when $k$ includes an asymptotic region, ð$\eth k$ is said to be stably marginal, where “marginal” refers to the vanishing of one of the null expansions, and “stably” refers to the absence of (in our language) any antinormal outward deformation of $k$. Here we omit the word “stably” because there is no other kind of “marginal.” In our discrete setting, when numerical values of the expansions are not invoked, the absence of antinormal outward deformations is the only natural way in which we are able define the notion of marginal. $k$ is future-marginally accessible from $a$ (past-marginally accessible from $a$) [^8] In the special case where a numerical expansion exists and $k$ contains an asymptotic region, ð$\eth k$ was called future- or past-minimar in Refs. Engelhardt:2018kcs ; Bousso:2019dxk .
$k$ is future- and past-marginal $k$ is a throat999ð$\eth k$ corresponds to a special case of a quantum extremal surface; it is minimal spatially and maximal in time. The other types of quantum extremal surfaces (bulge and bounce) Engelhardt:2023bpv do not have natural discrete definitions. $k$ is a throat accessible from $a$101010ð$\eth k$ was called outer-minimal in Ref. Engelhardt:2023bpv .
@@ 5.1 Antinormal Wedges
ð$\eth c$$\subset\eth a\cap b^{\prime}\sqcup H^{+}(a^{\prime})\cap H^{-}(b^{\prime})$
ð$\,\sqcup\eth b\cap a^{\prime}\sqcup H^{-}(a^{\prime})\cap H^{+}(b^{\prime})$
ð$\,\sqcup\eth a\cap\eth b~{}.$ (38)
We also note that ð$\eth\tilde{c}\subset J^{+}(\eth c)$, and that the portions of ð$\eth\tilde{c}$ that lie in the causal future of the first, second, and third line sets lie, respectively, on $H^{+}(a^{\prime})$, on $H^{+}(b^{\prime})$, and on $H^{+}(a)\cap H^{+}(b)$.
@@
The result follows immediately from Definitions 7 and 45, and the preceding Theorem. ∎
@@
Let $a$, $b$, and $h$ be wedges such that $h\subset a\cap b$. Suppose that $a$ is antinormal at all points on ð$\eth a\setminus\eth h$, and that $b$ is antinormal at all points on ð$\eth b\setminus\eth h$. Then $c=a\Cup b$ is antinormal at all points on ð$\eth c\setminus\eth h$.
@@
The proof is a straightforward generalization of the proof of the preceding Theorem. Instead of Eq. (39), we now define $\tilde{c}_{a}=(a\cup(\eth\tilde{c}\cap H^{+}(a^{\prime})\setminus H^{+}(h^{\prime})))^{\prime\prime}$, and similarly for $c_{a}$, $\tilde{c}_{b}$, and $c_{b}$. This ensures that only antinormal pieces of ð$\eth a$ and ð$\eth b$ are used when the QFC is invoked. It remains true that $c=c_{a}\Cup c_{b}$ and $\tilde{c}=\tilde{c}_{a}\Cup\tilde{c}_{b}$ since any portion of ð$\eth c$ or ð$\eth\tilde{c}$ that lies on $H^{+}(h^{\prime})$ must also lie on ð$\eth a$ or on ð$\eth b$ itself. ∎
@@ 5.2 Marginal Wedges and Throats
A wedge $a$ is said to be future-marginal at $p\in\eth a$ if
  1. 1.
    $a$ is future-nonexpanding at $p$.
  2. 2.
    There exists an open set $O\subset M$ containing $p$ such that no past-directed outward null deformation of $a$ away from $p$ within $O$ is future-nonexpanding. That is, no wedge $b\supset a$ with ð$\eth b\subset\eth a\cup[H^{-}(a^{\prime})\cap O]$, $p\notin\eth b$, is future-nonexpanding on all points in ð$\eth b\setminus\eth a$.
If this holds for all $p\in\eth a$ then $a$ is called future-marginal. Past-marginal is defined analogously.
@@
A wedge $a$ is said to be a throat at $p\in\eth a$ if $a$ is future- and past-marginal at $p$. If this holds for all $p\in\eth a$ then $a$ is called a throat.
@@
If the wedge $a$ is future-marginal at $p\in\eth a$, and if $\Theta_{\rm max}^{+}(a;p)$ exists, then $\Theta_{\rm max}^{+}(a;p)=0$.
@@
Since $a$ is future-nonexpanding, Theorem 41 implies that $\Theta_{\rm max}^{+}(a;p)\leq 0$. Suppose for contradiction that $\Theta_{\rm max}^{+}(a;p)<0$. By Def. 39, $\Theta_{\rm max}^{+}$ is continuous on its open domain, so there exists an open neighborhood $O$ of $p$ in $M$ such that $\Theta_{\rm max}^{+}(a;q)<0$ for all $q\in O\cap\eth a$. We may choose $O$ such that ð$\eth a\cap O\subset\eth a^{+}$. Assuming continuity of $\Theta_{\rm max}^{+}(a)$, we can smoothly deform $a$ outward along $H^{+}(a^{\prime})\cap O$, to obtain a wedge $b\supset a$ with $\Theta_{\rm max}^{+}(b;r)<0$ for all $r\in\eth b\setminus\eth a$. By Theorem 42, $b$ is a future-nonexpanding past outward deformation of $a$ at $p$. This contradicts the assumption that $a$ is future-marginal. ∎
@@
If the wedge $a$ is a throat at $p\in\eth a$, and if $\Theta_{\rm max}^{\pm}(a;p)$ both exist, then $\Theta_{\rm max}^{\pm}(a;p)=0$.

### v3

@@ 3.3 Special Case: Compressible States
We stress that in such settings, our novel definition of future-nonexpanding, Def. 28, and our Discrete Max-Focusing Conjecture, Conj. 33, are both still needed. Their purpose is to deal with edge points where the classical expansion is not well-defined, and this problem is unrelated to the quantum corrections to the area term.
@@ 3.4 Special Case: Numerical Max-Expansion
Let $a$ be a wedge whose edge $\eth a$ is acausal and $C^{2}$-smooth. (That is, no two points in $\eth a$ are causally connected; and $\eth a$ is a $C^{2}$ embedded submanifold.) Let $y$ collectively denote the transverse coordinates that label points in $\eth a$, and let $h$ be the determinant of the induced metric on $\eth a$. Let $V:\eth a\to[0,1]$, $y\to V(y)$ be a piecewise continuous function. Let $a^{+}_{V,\lambda}$ be the outward deformation of $a$ along the unique null generator of $H^{+}(a^{\prime})$ that originates at each point $y$ in $O$, by an affine distance $\lambda V(y)$. (Note that $a^{+}_{V,\lambda}$ exists for all sufficiently small $\lambda$.) Then there exists a continuous function $\Theta_{\rm max}^{+}:O\to\mathbb{R}$ such that
@@
The above definition can be generalized somewhat, though we will not try to do this carefully here: if $\eth a$ is not acausal or not smooth (or neither), $\Theta^{\pm}(a,p)$ can still be defined at any point $p\in\eth a$ that is acausal to the rest of $\eth a$ and possesses an open neighborhood $O$ such that $\eth a\cap O$ is a smooth embedded submanifold. Moreover, even at a point $p$ where $\eth a$ is not $C^{2}$, one of the expansions may exist: namely, if one of $H^{\pm}(a^{\prime})$ is smooth in a neighborhood of $p$.
@@
Since $a$ is future-nonexpanding at $p$, there exists an open neighborhood $O$ of $p$ such that any future-outward deformation $b$ along $H^{+}(a)$ that remains within $O$ obeys $H_{\rm max,gen}(b|a)\leq 0$. Let $V$ and $a^{+}_{V,\lambda}$ be defined as in Def. 40. Since $H_{\rm max,gen}(a^{+}_{V,\lambda}|a)$ is nonpositive for sufficiently small $\lambda$, and since Eq. (30) holds for all nonpositive functions $V:O\to\mathbb{R}_{0}^{+}$, one finds that $\Theta_{\rm max}^{+}\leq 0$. ∎
@@
By Def. 40, there exists an open neighborhood $O$ of $p$ such that $\Theta_{\rm max}^{+}(a;q)<0$ for all $q\in O$. We may now choose any piecewise continuous function $V:0\to[0,1]$, giving rise to a family of deformations $a^{+}_{V,\lambda}$ as described in Def. 40. Since $\Theta_{\rm max}^{+}(a;p)$ is strictly negative, by continuity, $H_{\rm max,gen}(a^{+}_{V,\lambda}|a)$ must be negative for all sufficiently small $\lambda$. After restricting $O$ to such values for all possible choices of $V$, Eq. (30) implies that Def. 28 is met. ∎
@@ 4.1 No Geodesics Enter Lightsheets
Discrete Max-Focusing would be violated already at the level of differential geometry if it was possible for new null geodesics to enter $L^{+}(a)$, i.e., if we could choose $\eth c$ to intersect null geodesics that were not present on $\eth b$. Then by choosing $\eth c$ to include only a small neighborhood of these geodesics and otherwise agree with $\eth b$, the area contribution to $H_{\rm max,gen}(c|b)$ would be expected to dominate and would be positive.
@@ 4.2 Classical Focusing
By Def. 28, $a$ is future-nonexpanding at any point $p$ where $\eth a$ is differentiable, if and only if the expansion $\theta^{+}$ of the null generators of $H^{+}(a^{\prime})$ is nonpositive at $a$. Then both Theorem 35 and Conjecture 33 follow immediately from the Raychaudhuri equation Wald:1984rg ,
@@ 4.3 Discrete Max-Focusing Implies the QNEC
Let $a$ be a wedge in Minkowski space (with $G=0$) whose edge $\eth a$ is acausal and smooth. Let $p\in\eth a$, and let the vector $k^{\mu}$ be null, future and outward-directed, and orthogonal to $\eth a$ at $p$. We require conditions on the extrinsic curvature of $\eth a$ at $p$ that are necessary for the QNEC to be regularization-scheme independent Akers:2017ttv . In particular, the classical null expansion and shear must vanish at $p$: $\theta^{+}=\varsigma^{+}_{\mu\nu}=0$. Let $\psi$ be the state of the quantum fields on $M$ (pure or mixed). Now consider $M$ with $n$ copies of the same quantum field theory, in the state $\psi^{\otimes n}$, and turn on $G\neq 0$. Our argument will involve taking the limit as $G$ is turned back off. Therefore, we can take $G$ to be arbitrarily small at the present step.
@@
Neither the identifications of $a(G,n)$ and $p(G,n)$, nor the deformation into $b(G,n)$ are unique, even while satisfying the criteria of Ref. Akers:2017ttv . The only important requirement is that they be chosen such that $p(G,n)\to p$, $a(G,n)\to a$, and $b(G,n)\to a$ as $G\to 0$ at fixed nn. [^4] To define an appropriate notion of continuity, one can represent the different spacetimes $M(G,n)$ as Minkowski space with a perturbative gravitational field. The latter property can be satisfied because $\Theta_{\rm max}^{+}[b(G,n);p(G,n)]-\theta^{+}[a(G,n);p(G,n)]\sim O(G)$.
We may now apply Discrete Max-Focusing to $b(G,n)$. Since $\Theta_{\rm max}^{+}[b(G,n);p(G,n)]<0$ and $b(G,n)$ can be chosen smooth, there exists an open neighborhood $O$ of $p(G,n)$ such that $\Theta_{\rm max}^{+}[b(G,n);q]<0$ for all $q\in\eth b(G,n)\cap O$; hence $\eth b(G,n)\cap O\subset\eth b(G,n)^{+}$. Let $c(G,n)\supset b(G,n)$ be an outward deformation of $b(G,n)$ along $H^{+}[b^{\prime}(G,n)]$ such that $\eth b(G,n)\setminus O=\eth c(G,n)\setminus O$. Let $q\in\eth c$ lie on the future null generator originating at $p$. We may choose $c(G,n)$ such that in the limit as $G\to 0$ at fixed $n$, it approaches a fixed wedge $c$ whose edge lies on $H^{+}(a^{\prime})$. (Recall that $a(G,n)\to a$, and $b(G,n)\to a$ as $G\to 0$ at fixed $n$.) Theorem 35 implies
@@
Here the quantum part is denoted with superscript $Q$ and defined as a functional derivative analogously to Eq. (30) (see Def. 40 for details):
@@
$\frac{1}{n}\left(\Theta_{\rm max}^{Q,+}[c;q]-\Theta_{\rm max}^{Q,+}[a;p]\right)\leq\frac{2\pi}{n\hbar}\int_{p}^{q}d\lambda\,\braket{T_{\mu\nu}}_{\psi^{\otimes n}}k^{\mu}k^{\nu}~.$ (36)
@@
$\Theta_{\rm vN}^{Q,+}[c;q]-\Theta_{\rm vN}^{Q,+}[a;p]\leq\frac{2\pi}{\hbar}\int_{p}^{q}d\lambda\,\braket{T_{\mu\nu}}_{\psi}k^{\mu}k^{\nu}~.$ (37)
@@ 5 Discrete-Max-Nonexpanding Wedges and Their Properties
$k$ is past- and future-nonexpanding at $p\in\eth k$ $k$ is antinormal at $p$ $k$ is accessible from aa [^6] This terminology was introduced in Ref. Akers:2023fqr for the special case where $a$ lies on the conformal boundary of Anti-de Sitter space.
$k$ is future- (past-) nonexpanding but no past (future) outward deformation of $k$ is future- (past-) nonexpanding $k$ is future-marginal (past-marginal) [^7] In the classical limit and when $k$ includes an asymptotic region, $\eth k$ is said to be stably marginal, where “marginal” refers to the vanishing of one of the null expansions, and “stably” refers to the absence of (in our language) any antinormal outward deformation of $k$. Here we omit the word “stably” because there is no other kind of “marginal.” In our discrete setting, when numerical values of the expansions are not invoked, the absence of antinormal outward deformations is the only natural way in which we are able define the notion of marginal. $k$ is future-marginally accessible from $a$ (past-marginally accessible from aa) [^8] In the special case where a numerical expansion exists and $k$ contains an asymptotic region, $\eth k$ was called future- or past-minimar in Refs. Engelhardt:2018kcs ; Bousso:2019dxk .
[…]
@@ 5.1 Antinormal Wedges
$\eth c$ $\subset\eth a\cap b^{\prime}\sqcup H^{+}(a^{\prime})\cap H^{-}(b^{\prime})$
$\,\sqcup\eth b\cap a^{\prime}\sqcup H^{-}(a^{\prime})\cap H^{+}(b^{\prime})$
$\,\sqcup\eth a\cap\eth b~.$ (38)
We also note that $\eth\tilde{c}\subset J^{+}(\eth c)$, and that the portions of $\eth\tilde{c}$ that lie in the causal future of the first, second, and third line sets lie, respectively, on $H^{+}(a^{\prime})$, on $H^{+}(b^{\prime})$, and on $H^{+}(a)\cap H^{+}(b)$.
@@
The result follows immediately from Definitions 7 and 46, and the preceding Theorem. ∎
@@
Let $a$, $b$, and $h$ be wedges such that $h\subset a\cap b$. Suppose that $a$ is antinormal at all points on $\eth a\setminus\eth h$, and that $b$ is antinormal at all points on $\eth b\setminus\eth h$. Then $c=a\Cup b$ is antinormal at all points on $\eth c\setminus\eth h$.
@@
The proof is a straightforward generalization of the proof of the preceding Theorem. Instead of Eq. (39), we now define $\tilde{c}_{a}=(a\cup(\eth\tilde{c}\cap H^{+}(a^{\prime})\setminus H^{+}(h^{\prime})))^{\prime\prime}$, and similarly for $c_{a}$, $\tilde{c}_{b}$, and $c_{b}$. This ensures that only antinormal pieces of $\eth a$ and $\eth b$ are used when the QFC is invoked. It remains true that $c=c_{a}\Cup c_{b}$ and $\tilde{c}=\tilde{c}_{a}\Cup\tilde{c}_{b}$ since any portion of $\eth c$ or $\eth\tilde{c}$ that lies on $H^{+}(h^{\prime})$ must also lie on $\eth a$ or on $\eth b$ itself. ∎
@@ 5.2 Marginal Wedges and Throats
A wedge $a$ is said to be future-marginal if
  1. I.
    $a$ is antinormal;
  2. II.
    $a$ is future-noncontracting.
Past-marginal is defined analogously.
@@
A wedge $a$ is said to be a throat if $a$ is future- and past-marginal.
@@
If the wedge $a$ is future-marginal, and if $\Theta_{\rm max}^{+}(a)$ exists, then $\Theta_{\rm max}^{+}(a)=0$.
@@
Since $a$ is future-nonexpanding, Theorem 42 implies that $\Theta_{\rm max}^{+}(a)\leq 0$. Suppose for contradiction that $\Theta_{\rm max}^{+}(a)<0$. By Def. 40, $\Theta_{\rm max}^{+}$ is continuous on its open domain, so there exists an open neighborhood $O$ of $\eth a$ in $M$ such that $\Theta_{\rm max}^{+}(a)<0$ for all $q\in O\cap\eth a$. Assuming continuity of $\Theta_{\rm max}^{+}(a)$, we can smoothly deform $a$ outward along $H^{+}(a^{\prime})\cap O$, to obtain a wedge $b\supset a$ with $\Theta_{\rm max}^{+}(b;r)<0$ for all $r\in\eth b\setminus\eth a$. By Theorem 43, $b$ is a future-nonexpanding past outward deformation of $a$. This contradicts the assumption that $a$ is future-marginal. ∎
@@
If the wedge $a$ is a throat, and if $\Theta_{\rm max}^{\pm}(a)$ both exist, then $\Theta_{\rm max}^{\pm}(a)=0$.